OpenStudy (anonymous):

How to find the intercepts and asymptotes of f(x)=2x^2+x-2/x^2-1

OpenStudy (anonymous):

Is the function\[f(x)=\frac{ 2x^2+x-2 }{ x^2-1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK lets start with intercepts. What is the interception with y-axis? Isnt it when x=0?

OpenStudy (anonymous):

sorry, intercept with y-axis

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so, what is the value of f(x) when x=0?

OpenStudy (anonymous):

f(0)=2(0)^2+0-2/(0)^2-1=-2/-1=2

OpenStudy (anonymous):

Perfect! Now The x-axis crossing! Arent they the values of x that make y=0? And arent the values that make y=0 the same values that make the numerator=0?

OpenStudy (anonymous):

so I have to get the numerator equal to zero

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

so 0=2x^2+0-2, how can I solve that

OpenStudy (anonymous):

i meant +x not +0

OpenStudy (anonymous):

You have to solve:\[2x^2+x-2=0\]the formula to solve an equation like:\[ax^2+bx+c=0\rightarrow x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]I dont believe you didnt know this formula

OpenStudy (anonymous):

in your case a=2, b=1 and c=-2, find the two solutions

OpenStudy (anonymous):

I get:x1=0.78 and x2=-1.28 do you ge the same?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yeah how did you get the .78 and 1.28 when you get left off with \[-1\pm \sqrt{5}\div4\]

OpenStudy (anonymous):

so, for you b^2-4ac = 5?

OpenStudy (anonymous):

a=2 b=1 c=-2-->b^2-4ac=1-4·2·(-2)=1+16=17

OpenStudy (anonymous):

oops i found my mistake , ok I'm with you now

OpenStudy (anonymous):

so its \[-1\pm \sqrt{17}\div4\]

OpenStudy (anonymous):

Those are the cross points for x-axis

OpenStudy (anonymous):

Now find the roots of the denominator. It is straightforward

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

what are the values of x that make the denomnator =0?

OpenStudy (anonymous):

I am sorry, I have to go. To be continued tomorrow, bye

OpenStudy (anonymous):

i've figured it out and the asymptotes. can you help me find the x-intercept of x^2-2x+3/x+2